OpenStudy (anonymous):

3SIN2THETA - 2SINTHEA = 0

OpenStudy (anonymous):

is it this problem???: \(\large 3sin^2\theta - 2sin\theta = 0 \)

OpenStudy (anonymous):

3sin2theta - 2sintheta

OpenStudy (anonymous):

NO exponents in the first part.

OpenStudy (anonymous):

\(\large 3sin(2\theta)-2sin\theta=0 \) like this?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok... use the double angle formula for sine...: \(\sin(2\theta) = 2sin\theta cos\theta \)

OpenStudy (anonymous):

replace that into your original equation... what you got now?

OpenStudy (anonymous):

\[\huge 3\color {red}{sin(2\theta)}-2sin\theta=0 \] \[\huge 3\color {red}{2sin\theta \cdot cos\theta}-2sin\theta=0 \]

OpenStudy (anonymous):

6sintheta(costheta) - 2sintheta =0

OpenStudy (anonymous):

yep.... now factor out a sin(theta)..... from the left side...

OpenStudy (anonymous):

sintheta (6costheta - 2) = 0

OpenStudy (anonymous):

great.... now just set each factor equal to zero and solve....

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

solve both: \(\large sin\theta = 0 \) and \(\large 6cos\theta - 2=0 \)

OpenStudy (anonymous):

are you looking for the general solution or solutions within [0, 2pi) ????

OpenStudy (anonymous):

solutions like Pi 2Pi

OpenStudy (anonymous):

so the general solution...????

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

from that fist equation, \(\sin\theta=0 \)---> \(\theta=n\cdot \pi \), where n is an integer....

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

from the second eqaution, \(\large 6 cos\theta - 2=0 \rightarrow cos\theta=\frac{1}{3} \) so \(\large \theta=Arccos\frac{1}{3} + \)

OpenStudy (anonymous):

Arc cos 1/3 +?